OpenStudy (anonymous):

32 to the power of -3/5

ganeshie8 (ganeshie8):

\(\huge 32 ^{\frac{-3}{5}} \\ = \huge 2^{5*\frac{-3}{5}} \\ = \huge 2^{\cancel{5}*\frac{-3}{\cancel{5}}} \\ \huge = 2^{-3} \\ \huge = \frac{1}{2^3} \\ \huge = \frac{1}{8} \)